OpenStudy (anonymous):

Solve for x: x = –4 x = 5 x = –4 and x = 5 No Solution

OpenStudy (anonymous):

\[\frac{ 1 }{2 }+\frac{ 4 }{ 2x }=\frac{ x+4 }{ 10 }\]